2K+AR = R-3Y Must solve for R
Answer by nyc_function(2741)

2K+AR = R-3Y Must solve for R
2K+AR = R-3Y
AR - R = -2K - 3Y
R(A - 1) = -2K - 3Y
R = (-2K - 3Y)/(A - 1)
Done!